OpenStudy (anonymous):

What is the slope of the line that passes through the points (-3, 5) and (1, 7)? a) 1/2 b)1 c)2

pooja195 (pooja195):

\[\huge~\rm~\frac{ y_2-y_1 }{ x_2-x_1 }\]

pooja195 (pooja195):

We need to find the slope first :) (x1,y1)(x2,y2)

OpenStudy (anonymous):

ok @pooja195

pooja195 (pooja195):

\[\huge~\rm~\frac{ 7-5 }{ 1+3}=slope \]

OpenStudy (anonymous):

thanks its 1/2 @pooja195

pooja195 (pooja195):

no its not.

OpenStudy (anonymous):

its not??

pooja195 (pooja195):

look at what i did above

pooja195 (pooja195):

\[\huge~\rm~\frac{ 7-5 }{ 1+3}=slope \]

pooja195 (pooja195):

^^^^^^^^^

OpenStudy (anonymous):

I did do that but I got 2/4 and I simplified

pooja195 (pooja195):

oh right oops sorry!

OpenStudy (anonymous):

its alright @pooja195

pooja195 (pooja195):

\[\huge~\rm~y-y_1=m(x-x_1)\] ok now we have our slope m=slope so 1/2 we can pick and point (1,7) will be easier

pooja195 (pooja195):

(x1,y1)